OpenStudy (simmy2497):

PLease help me!!!! Write the point-slope form of the equation of the line that passes through the origin and has a slope of 2. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (templeguy):

answer mine and i'll answer urs

OpenStudy (simmy2497):

whats ur question??

OpenStudy (templeguy):

I'll send it in message

OpenStudy (templeguy):

here's the answer: y-0 = 2 (x-0)y=2x